OpenStudy (anonymous):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

what do you think it is @bubblegumcrazy

OpenStudy (thefurball):

you need to find the slope of the line AB. Since the line you are finding is parallel to AB, the slopes will be the same. you can then use the slope and the point to write the new equation in point-slope form.

OpenStudy (anonymous):

y = 5x - 15 y = 1/5x -29/5 y = 5x + 15 y = 1/5x +21/5

OpenStudy (anonymous):

uuuummmm thinking:)

OpenStudy (thefurball):

ill give you a hint, AB is in y = mx + b (m is the slope) point slope form is \[y - y _{1} = m(x - x _{1})\] you can use point slope form and solve for y to get it in slope-intercept form

OpenStudy (anonymous):

I'm really bad at these ones,....ummm could it be C?

OpenStudy (thefurball):

ok, so the slope is 5 because that's the slope in AB. the new point is (4,5). Plug that in to point slope to get \[y - 5 = 5(x -4)\] then, solving for y, we get \[y = 5x - 20 +5 = 5x - 15\]C was close though! If you need further help, message me!

OpenStudy (anonymous):

oh then it must be A